OpenStudy (anonymous):

2x times square root of x?

OpenStudy (anonymous):

are you trying to write the expression?

OpenStudy (anonymous):

yeah dude (2x+3)(square root of x -6)

OpenStudy (anonymous):

please help asap pls

OpenStudy (anonymous):

is it \[\sqrt{x} ~ or~ \sqrt{x-6}\]

OpenStudy (anonymous):

is square root of x -6

OpenStudy (anonymous):

the first one minus 6

OpenStudy (anonymous):

\[(2x+3)(\sqrt{x}-6 \[= 2x \sqrt{x} -12x + 3\sqrt{x} - 18\] Here i used the multiplicative property